OpenStudy (anonymous):

solve for y 2x + 3y =18 (write in slope-intercept form)

OpenStudy (anonymous):

\[2x+3y=18\] \[3y=-2x+18\]\[3y \div 3 =- 2x \div 3 + 18 \div 3\]

OpenStudy (anonymous):

y=-2/3x+6